# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

### v1

@@ 1.1. Local reconstruction analysis (LRA)
Let $\mathcal{R}$ be the generalized Radon transform (GRT); $f(x)$, $x\in\mathcal{U}\subset\mathbb{R}^{n}$, be the function to be reconstructed; and $g(y)$, $y\in\mathcal{V}\subset\mathbb{R}^{n}$, where $g=\mathcal{R}f$, be the GRT of $f$. The open sets $\mathcal{U}$ and $\mathcal{V}$ represent the image and data domains, respectively. The values of $g(y)$ are given on a rectangular grid $y_{j}$, $j\in\mathbb{Z}^{n}$. The step size of the grid along each coordinate is proportional to some small parameter $\epsilon>0$.
Consider a reconstruction formula of the form $f_{0}=\mathcal{R}^{*}\mathcal{B}g$, where $\mathcal{R}$ is the adjoint transform and $\mathcal{B}$ is a Pseudo-Differential Operator ($\Psi$DO). Depending on the choice of $\mathcal{B}$, the reconstruction can be theoretically exact ($f_{0}=f$), quasi-exact ($f_{0}-f$ is less singular than $f$), edge-enhancing ($f_{0}$ is more singular than $f$), and smoothing of a finite degree ($f_{0}$ is less singular than $f$).
Next, let $\varphi$ be a compactly supported, sufficiently smooth interpolation kernel, and let $g_{\epsilon}(y)$, $y\in\mathcal{V}$, denote the function obtained by interpolating the values $g(y_{j})$. Denote $f_{\epsilon}=\mathcal{R}^{*}\mathcal{B}g_{e}$. Thus, $f_{\epsilon}$ is reconstructed from discrete data.
LRA is a powerful and flexible technique proposed by the author to study images $f_{\epsilon}$ reconstructed from discrete GRT data. The main idea of LRA is to obtain a simple formula to accurately approximate $f_{\epsilon}$ in an $\epsilon$-neighborhood of a point, $x_{0}\in\mathcal{U}$. To illustrate the main idea of LRA consider an example. Let $f$ be a piece-wise smooth real-valued function with singularities along some hypersurface $\mathcal{S}$. More precisely, $f$ is conormal with respect to $\mathcal{S}$ [21, Definition 18.2.6]. Fix any $A>0$. We show that, in a variety of settings, the following limit exists
@@
To illustrate one such setting, suppose $f$ has a jump across $\mathcal{S}$, where $\mathcal{S}\subset\mathbb{R}^{2}$ is a smooth curve. Suppose also that the reconstruction is either exact or quasi-exact. In this case $\mathcal{R}^{*}\mathcal{B}\mathcal{R}$ is a $\Psi$DO with the principal symbol 1 and $\nu=0$. Pick $x_{0}\in\mathcal{S}$. Under some mild conditions on $\mathcal{S}$ and $x_{0}$, we show
(1.2) $\lim_{\epsilon\to 0}f_{\epsilon}(x_{0}+\epsilon\check{x})=f_{0}(x_{0}^{-})+\Delta f(x_{0})\text{DTB}(\check{x};x_{0}),\ |\check{x}|<A,$
@@
    $f_{0}(x_{0}^{-})$ is the value of the reconstruction from continuous data, $f_{0}$, on one side of $\mathcal{S}$,
@@
    $\Delta f(x_{0}):=f_{0}(x_{0}^{+})-f_{0}(x_{0}^{-})$ is the value of the jump of $f$ across $\mathcal{S}$ at $x_{0}$, and
@@
The DTB function depends only on the curvature of $\mathcal{S}$ at $x_{0}$ and on the interpolation kernel, $\varphi$. When $\epsilon$ is sufficiently small, the right-hand side of (1.2) is an accurate approximation of $f_{\epsilon}$, and the DTB function accurately describes the smoothing of the jump of $f$ in $f_{\epsilon}$ due to the discrete nature of the tomographic data.
@@
Even though LRA applies in a wide variety of settings, one of its major limitations is that so far it has only been applied when $f_{\epsilon}$ is computed by a linear reconstruction algorithm, i.e., by an application of a reconstruction formula to data: $f_{\epsilon}=\mathcal{R}^{*}\mathcal{B}g_{\epsilon}$. Very few GRTs admit an exact inversion formula. In such cases, at best, formula-based reconstruction can only guarantee an accurate recovery of the singularities of $f$ (which we called quasi-exact reconstruction above).
@@ 1.2. New result - LRA for iterative reconstruction
(1.3) $f_{\epsilon}=\text{argmin}_{f\in H_{c}^{1}(\mathcal{U}_{b})}\Psi(f),\ \Psi(f):=\|\mathcal{R}f-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+\kappa\epsilon^{3}\|\nabla f\|_{L^{2}(\mathcal{U})}^{2},$
where $\kappa>0$ is the regularization parameter (which is predetermined and independent of $\epsilon$). In the spirit of LRA, the factor $\epsilon^{3}$ in front of the Tikhonov regularization term ensures that the resolution of the reconstruction is at the native scale $\epsilon$.
To solve (1.3) numerically, one usually assumes that $f_{\epsilon}(x)=\sum_{l}c_{l}\psi_{l}(x)$, where $\psi_{l}$ are some basis functions, and their number is of the same order of magnitude as the number of data points. In this case, the reconstruction consists of computing the coefficients $c_{l}$. For conventional analysis of reconstruction at a scale $\gg\epsilon$, the effects of this finite-dimensional approximation of $f_{\epsilon}$ are largely invisible. However, LRA is designed to study reconstruction at the native scale, $\epsilon$. This is the scale of greatest practical importance. At this scale, the two types of effects, (a) due to the discrete nature of data and (b) due to the representation of $f_{\epsilon}$ as a linear combination of basis functions, become comparable and relevant.
@@
An important feature of the minimization problem (1.3) is that it looks for a solution, $f_{\epsilon}$, in an (infinite dimensional) Hilbert space. Of course, this is numerically impossible. In the end, we still have to find a solution in a finite-dimensional vector space. However, numerically, the effects of this replacement of the solution space can be made as small as one likes relative to the effects of discrete data. We consider (1.3) because our goal is to study the most fundamental limitation on image quality, namely the limitation due to discrete data. While the effects of finite-dimensional approximation of a solution are less fundamental (because they can be effectively mitigated), they are important from the point of view of computational complexity. Nevertheless, this line of research is beyond the scope of the paper and will be the subject of future work.
@@ 1.3. Organization of the paper
The rest of the paper contains the proofs of the two theorems. In section 4 we obtain a formula for $\mathcal{R}f$ in a neighborhood of its singular support, (4.13). In section 5 we obtain an alternative formula for computing $f_{\epsilon}$, equation (5.6). The equation is of the form $\mathcal{T}f_{\epsilon}=F_{\epsilon}$, where $\mathcal{T}$ is some $\Psi$DO, and $F_{\epsilon}$ is computed from the data $g(y_{j})$. Also we find another function, $G_{\epsilon}$, such that (a) $G_{\epsilon}$ is easier to analyze than $f_{\epsilon}$ and (b) the local behavior of $f_{\epsilon}$ and $G_{\epsilon}$ near $x_{0}$ are the same from the perspective of LRA, see Lemma 5.2. In section 6 we prove that $\|F_{\epsilon}\|_{L^{\infty}(\mathcal{U})}$ is uniformly bounded when $0<\epsilon\ll 1$. In section 7 we compute the limit
@@
where $F_{\epsilon}^{(l)}(x)$ is the leading singular term of $F_{\epsilon}(x)$ near $x_{0}$, see (7.18). In section 8 we compute the DTB function, $\Delta G_{0}$, as the limit
(1.5) $\Delta G_{0}(\check{x}):=\lim_{\epsilon\to 0}\big{(}G_{\epsilon}(x_{0}+\epsilon\check{x})-G_{\epsilon}(x_{0})\big{)},$
see equation (LABEL:rhs_lim_5). This calculation assumes that the singularity of $f$ at $x_{0}$ is visible from the data only once. This completes the proof of Theorem 3.6.
@@ 2. Basic notation, function spaces, symbols, operators
Let $f$ and $g$ be two functions defined on a domain $U$. We say $f(x)\asymp g(x)$ for $x\in U$ if there exist $c_{1,2}>0$ such that
(2.1) $c_{1}\leq f(x)/g(x)\leq c_{2}\text{ if }g(x)\not=0\text{ and }f(x)=0\text{ if }g(x)=0$
for any $x\in U$.
@@
Another type is the Hölder spaces $C^{s}(\mathbb{R}^{n})$, $s>0$, $s\not\in\mathbb{N}$. If $s=k+\gamma$, $k\in\mathbb{N}_{0}$, $10<\gamma<$ 1, then $C^{s}(\mathbb{R}^{n})$ is the space of $C^{k}(\mathbb{R}^{n})$ functions (or, $L^{\infty}(\mathbb{R}^{n})$ functions if $k=0$), which have Hölder continuous $k$-th order derivatives (see [4, Definition 5.28]), with the norm
(2.3) $\begin{split}\|f\|_{C^{s}}:=&\|f\|_{C^{k}}+\max_{|m|=k}\sup_{x\in\mathbb{R}^{n},|h|>0}\frac{|f^{(m)}(x+h)-f^{(m)}(x)|}{|h|^{\gamma}},\ k>0,\\ \|f\|_{C^{s}}:=&\|f\|_{L^{\infty}}+\sup_{x\in\mathbb{R}^{n},|h|>0}\frac{|f(x+h)-f(x)|}{|h|^{\gamma}},\ k=0.\end{split}$
@@
For $s>0$, $s\not\in\mathbb{N}$, Hölder spaces coincide with Zygmund spaces, $C_{*}^{s}(\mathbb{R}^{n})$. To define the latter, pick any $\mu_{0}\in C_{0}^{\infty}(\mathbb{R}^{n})$ such that $\mu_{0}(\eta)=1$ for $|\eta|\leq 1$, $\mu_{0}(\eta)=0$ for $|\eta|\geq 2$. Define $\mu_{j}(\eta):=\mu_{0}(2^{-j}\eta)-\mu_{0}(2^{-j+1}\eta)$, $j\in\mathbb{N}$ [1, Section 5.4]. Then
@@
which extends to tempered distributions [20, section 7.1]. We have $C_{*}^{s}(\mathbb{R}^{n})=C^{s}(\mathbb{R}^{n})$ and the norms $\|\cdot\|_{C_{*}^{s}(\mathbb{R}^{n})}$, $\|\cdot\|_{C^{s}(\mathbb{R}^{n})}$ are equivalent, for any $s>0$, $s\not\in\mathbb{N}$ [4, Remark 5.31].
@@
$H_{c}^{s}(U)$ is the collection of all $f\in H^{s}(\mathbb{R}^{n})$ such that $f\equiv 0$ on $\mathbb{R}^{n}\setminus K$ for some compact $K\subset U$. $H_{loc}^{s}(U)$ is the space of all distributions on $U$ such that $\chi f\in H^{s}(U)$ for any $\chi\in C_{0}^{\infty}(U)$. See [52, section 5, Notation and Background]
@@
(2.7) $\begin{split}&|\partial_{x}^{m}\tilde{B}(x,\xi)|\leq c_{m}\ \forall m\in\mathbb{N}_{0}^{n},x\in U,0<|\xi|\leq 1;\\ &|\partial_{x}^{m_{1}}\partial_{\xi}^{m_{2}}\tilde{B}(x,\xi)|\leq c_{m_{1},m_{2}}|\xi|^{r-|m_{2}|},\ \forall m_{1}\in\mathbb{N}_{0}^{n},m_{2}\in\mathbb{N}_{0}^{N},x\in U,|\xi|\geq 1;\end{split}$
for some constants $c_{m},c_{m_{1},m_{2}}>0$.
@@
is denoted $L^{r}(U)$. As is known, $L^{r}(U):H_{c}^{s}(U)\to H_{loc}^{s-r}(U)$ is continuous [52, Theorem 2.1, Section I.2].
@@
be a nondegenerate phase function [16, Definition 11.1]. The latter property means that
@@
    The differentials $\text{d}_{(x,y,\Theta)}(\partial\Phi/\partial\Theta_{l})$, $1\leq l\leq N$, are linearly independent on the set $\Sigma$:
    (2.10) $\Sigma:=\{(x,y,\Theta)\in U\times V\times(\mathbb{R}^{N}\setminus 0_{N}):\ \Phi_{\Theta}^{\prime}(x,y,\Theta)=0_{N}\}.$
The phase $\Phi$ determines a map $T^{*}U\to T^{*}V$ with the graph
(2.11) $\mathcal{C}:=\{(y,\text{d}_{y}\Phi(x,y,\Theta));(x,-\text{d}_{x}\Phi(x,y,\Theta)):(x,y,\Theta)\in\Sigma\}.$
$I^{r}(V\times U\times\mathcal{C})$ denotes the vector space of FIOs of order $r$ given by [16, section 11]

### v2

@@ 1.1. Local reconstruction analysis (LRA)
Let $\mathcal{R}$ be the generalized Radon transform (GRT), which integrates a function $f$ on $\mathcal{U}\subset\mathbb{R}^{n}$ over a family of surfaces parameterized by a set $\mathcal{V}\subset\mathbb{R}^{n}$. Thus, $f$ is the function to be reconstructed. Let $\hat{f}(y)=(\mathcal{R}f)(y)$, $y\in\mathcal{V}$, denote the GRT of $f$. The open sets $\mathcal{U}$ and $\mathcal{V}$ represent the image and data domains, respectively. The values of $\hat{f}(y)$ are given on a rectangular grid $y_{j}$, $j\in\mathbb{Z}^{n}$. The step size of the grid along each coordinate is proportional to a small parameter $\epsilon>0$.
Consider a reconstruction formula of the form $f_{0}=\mathcal{R}^{*}\mathcal{B}\hat{f}$, where $\mathcal{R}$ is the adjoint transform and $\mathcal{B}$ is a Pseudo-Differential Operator ($\Psi$DO). Depending on the choice of $\mathcal{B}$, the reconstruction can be theoretically exact ($f_{0}=f$), quasi-exact ($f_{0}-f$ is less singular than $f$), edge-enhancing ($f_{0}$ is more singular than $f$), or smoothing of a finite degree ($f_{0}$ is less singular than $f$). The common property of each of these kinds of reconstructions is that, microlocally, $\text{WF}(f)=\text{WF}(f_{0})$.
Next, let $\varphi$ be a compactly supported, sufficiently smooth interpolation kernel, and let $\hat{f}_{\epsilon}(y)$, $y\in\mathcal{V}$, denote the function obtained by interpolating the values $\hat{f}(y_{j})$, see eq. (3.14). Denote $f_{\epsilon}=\mathcal{R}^{*}\mathcal{B}\hat{f}_{\epsilon}$. Thus, $f_{\epsilon}$ is reconstructed from discrete data.
LRA is a powerful and flexible technique proposed by the author to study images $f_{\epsilon}$ reconstructed from discrete GRT data. The reconstruction can be of any kind, e.g., exact, quasi-exact, edge-enhancing, or smoothing as described above. The main idea of LRA is to obtain a simple formula to accurately approximate $f_{\epsilon}$ in an $\epsilon$-neighborhood of a point, $x_{0}\in\mathcal{U}$. We assume that $f$ is conormal with respect to some hypersurface $\mathcal{S}$ [24, Definition 18.2.6]. Fix any $x_{0}\in\mathcal{S}$ and $A>0$. We show that, in a variety of settings, the following limit exists
@@
To help understand (1.1), we consider an example of a GRT in the plane. Suppose $f$ has a jump across $\mathcal{S}$, where $\mathcal{S}\subset\mathbb{R}^{2}$ is a smooth curve. Suppose also that the reconstruction is either exact or quasi-exact. In this case $\mathcal{R}^{*}\mathcal{B}\mathcal{R}$ is a $\Psi$DO with the principal symbol 1 and $\nu=0$. Let $\vec{\Theta}_{0}$ be a unit vector perpendicular to $\mathcal{S}$ at $x_{0}$. Under some mild conditions on $\mathcal{S}$ and $x_{0}$, we show
(1.2) $\lim_{\epsilon\to 0}f_{\epsilon}(x_{0}+\epsilon\check{x})=f_{0}(x_{0}^{-})+\Delta f(x_{0})\text{DTB}(\vec{\Theta}_{0}\cdot\check{x};x_{0}),\ |\check{x}|<A,$
@@
    $f_{0}(x_{0}^{\pm}):=\lim_{t\to 0^{\pm}}f_{0}(x_{0}+t\vec{\Theta}_{0})$;
@@
    $\Delta f(x_{0}):=f_{0}(x_{0}^{+})-f_{0}(x_{0}^{-})$ is the value of the jump of $f$ across $\mathcal{S}$ at $x_{0}$; and
@@
The DTB function depends only on the interpolation kernel, $\varphi$. When $\epsilon$ is sufficiently small, the right-hand side of (1.2) is an accurate approximation of $f_{\epsilon}$, and the DTB function accurately describes the smoothing of the jump of $f$ in $f_{\epsilon}$ due to the discrete nature of the tomographic data.
@@
Even though LRA applies in a wide variety of settings, one of its major limitations is that so far it has only been applied when $f_{\epsilon}$ is computed by a linear reconstruction algorithm, i.e., by an application of a reconstruction formula to data: $f_{\epsilon}=\mathcal{R}^{*}\mathcal{B}\hat{f}_{\epsilon}$. Very few GRTs admit an exact inversion formula. In such cases, at best, formula-based reconstruction can only guarantee an accurate recovery of the singularities of $f$ (which we called quasi-exact reconstruction above).
@@ 1.2. New result - LRA for iterative reconstruction
(1.3) $f_{\epsilon}=\text{argmin}_{f\in H_{0}^{1}(\mathcal{U}_{b})}\Psi(f),\ \Psi(f):=\|\mathcal{R}f-\hat{f}_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+\kappa\epsilon^{3}\|\partial_{x}f\|_{L^{2}(\mathcal{U})}^{2},$
where $\kappa>0$ is the regularization parameter (which is predetermined and independent of $\epsilon$), $U_{b}\Subset U$ is a domain, and $H_{0}^{1}(\mathcal{U}_{b})$ is the closure of $C_{0}^{\infty}(U_{b})$ in the $H^{1}$ norm. In the spirit of LRA, the factor $\epsilon^{3}$ in front of the Tikhonov regularization term ensures that the resolution of the reconstruction is at the native scale $\epsilon$.
To solve (1.3) numerically, one usually assumes that $f_{\epsilon}(x)=\sum_{l}c_{l}\psi_{l}(x)$, where $\psi_{l}$ are some basis functions, and their number is of the same order of magnitude as the number of data points. The reconstruction consists of computing the coefficients $c_{l}$. For conventional analysis of reconstruction at a scale $\gg\epsilon$, the effects of this finite-dimensional approximation of $f_{\epsilon}$ are largely invisible. However, LRA is designed to study reconstruction at the native scale, $\epsilon$. This is the scale of greatest practical importance. At this scale, the two types of effects, (a) due to the discrete nature of data and (b) due to the representation of $f_{\epsilon}$ as a linear combination of basis functions, become comparable and relevant.
@@
An important feature of the minimization problem (1.3) is that it looks for a solution, $f_{\epsilon}$, in an (infinite dimensional) Hilbert space. Of course, this is numerically impossible. In the end, we still have to find a solution in a finite-dimensional vector space. However, numerically, the effects of this replacement of the solution space can be made as small as one likes relative to the effects of discrete data. We consider (1.3) because our goal is to study the most fundamental limitation on image quality, namely the limitation due to discrete data. While the effects of finite-dimensional approximation of a solution are less fundamental (because they can be effectively mitigated), they are important from the point of view of computational complexity. Nevertheless, this line of research is beyond the scope of the paper.
@@ 1.3. Organization of the paper
The rest of the paper contains the proofs of the two theorems. In Section 4 we obtain a formula for $\mathcal{R}f$ in a neighborhood of its singular support, (4.13). In Section 5 we obtain an alternative formula for computing $f_{\epsilon}$, equation (5.6). The equation is of the form $\mathcal{T}f_{\epsilon}=F_{\epsilon}$, where $\mathcal{T}$ is some $\Psi$DO, and $F_{\epsilon}$ is computed from the data $\hat{f}(y_{j})$. Also, we find another function, $\mathcal{G}_{\epsilon}$, such that (a) $\mathcal{G}_{\epsilon}$ is easier to analyze than $f_{\epsilon}$ and (b) the local behavior of $f_{\epsilon}$ and $\mathcal{G}_{\epsilon}$ near $x_{0}$ are the same from the perspective of LRA, see Lemma 5.2. In Section 6 we prove that $\|F_{\epsilon}\|_{L^{\infty}(\mathcal{U})}$ is uniformly bounded when $0<\epsilon\ll 1$. In Section 7 we compute the limit
@@
where $F_{\epsilon}^{(l)}(x)$ is the leading singular term of $F_{\epsilon}(x)$ near $x_{0}$, see (7.20). In Section 8 we compute the DTB function, $\Delta\mathcal{G}_{0}$, as the limit
(1.5) $\Delta\mathcal{G}_{0}(\check{x}):=\lim_{\epsilon\to 0}\big{(}\mathcal{G}_{\epsilon}(x_{0}+\epsilon\check{x})-\mathcal{G}_{\epsilon}(x_{0})\big{)},$
see equation (LABEL:rhs_lim_5). The limits in (1.4) and (1.5) are uniform in $\check{x}$ confined to any bounded set. The calculation assumes that the singularity of $f$ at $x_{0}$ is visible from the data only once. This means that there is only one $y\in\mathcal{V}$ such that $\mathcal{S}_{y}$ is tangent to $\mathcal{S}$ at $x_{0}$. This completes the proof of Theorem 3.6.
@@ 2. Basic notation, function spaces, symbols, operators
Let $f$ and $g$ be two functions defined on a domain $U$. We say $f(x)\asymp g(x)$ for $x\in U$ if there exist $c_{1,2}>0$ such that $f(x)\leq c_{1}g(x)$ and $g(x)\leq c_{2}f(x)$ for any $x\in U$.
We identify tangent and cotangent spaces on Euclidean spaces with the underlying Euclidean spaces. Thus, for a function $f(x)$ defined on $\mathbb{R}^{n}$ we have
(2.1) $\begin{split}&\partial_{x}f(x):=f_{x}^{\prime}(x):=(\partial_{x_{1}}f(x),\dots,\partial_{x_{n}}f(x))^{T},\\ &\text{d}_{x}f=(\partial_{x_{1}}f(x),\dots,\partial_{x_{n}}f(x)),\\ &|\partial_{x}f(x)|=|\text{d}_{x}f(x)|=\big{[}(\partial_{x_{1}}f(x))^{2}+\dots+(\partial_{x_{n}}f(x))^{2}\big{]}^{1/2}.\end{split}$
We use $f_{x}^{\prime}$ or $\partial_{x}f$ interchangeably depending on what is more convenient from the notational perspective.
@@
Another type is the Hölder spaces $C^{s}(\mathbb{R}^{n})$, $s>0$, $s\not\in\mathbb{N}$. If $s=k+\gamma$, $k\in\mathbb{N}_{0}$, $10<\gamma<$ 1, then $C^{s}(\mathbb{R}^{n})$ is the space of $C^{k}(\mathbb{R}^{n})$ functions (or, $L^{\infty}(\mathbb{R}^{n})$ functions if $k=0$), which have Hölder continuous $k$-th order derivatives (see [5, Definition 5.28]), with the norm
(2.3) $\begin{split}\|f\|_{C^{s}}:=&\|f\|_{C^{k}}+\max_{|m|=k}\sup_{x\in\mathbb{R}^{n},|h|>0}\frac{|\partial_{x}^{m}(f(x+h)-f(x))|}{|h|^{\gamma}},\ k>0,\\ \|f\|_{C^{s}}:=&\|f\|_{L^{\infty}}+\sup_{x\in\mathbb{R}^{n},|h|>0}\frac{|f(x+h)-f(x)|}{|h|^{\gamma}},\ k=0.\end{split}$
@@
For $s>0$, $s\not\in\mathbb{N}$, Hölder spaces coincide with Zygmund spaces, $C_{*}^{s}(\mathbb{R}^{n})$. To define the latter, pick any $\mu_{0}\in C_{0}^{\infty}(\mathbb{R}^{n})$ such that $\mu_{0}(\eta)=1$ for $|\eta|\leq 1$, $\mu_{0}(\eta)=0$ for $|\eta|\geq 2$. Define $\mu_{j}(\eta):=\mu_{0}(2^{-j}\eta)-\mu_{0}(2^{-j+1}\eta)$, $j\in\mathbb{N}$ [5, Definition 5.30] and [1, Section 6.2]. Then
@@
which extends to tempered distributions [25, Section 7.1]. We have $C_{*}^{s}(\mathbb{R}^{n})=C^{s}(\mathbb{R}^{n})$ and the norms $\|\cdot\|_{C_{*}^{s}(\mathbb{R}^{n})}$, $\|\cdot\|_{C^{s}(\mathbb{R}^{n})}$ are equivalent, for any $s>0$, $s\not\in\mathbb{N}$ [5, Remark 5.31].
@@
$H_{0}^{s}(U)$ is the closure of $C_{0}^{\infty}(U)$ in $H^{s}(\mathbb{R}^{n})$. $H_{loc}^{s}(U)$ is the space of all distributions on $U$ such that $\chi f\in H^{s}(\mathbb{R}^{n})$ for any $\chi\in C_{0}^{\infty}(U)$. See [55, Section 5, Notation and Background]
@@
(2.7) $\begin{split}&|\partial_{x}^{m}\tilde{B}(x,\xi)|\leq c_{m}|\xi|^{-n+\delta}\ \forall m\in\mathbb{N}_{0}^{n},x\in U,0<|\xi|\leq 1;\\ &|\partial_{x}^{m_{1}}\partial_{\xi}^{m_{2}}\tilde{B}(x,\xi)|\leq c_{m_{1},m_{2}}|\xi|^{r-|m_{2}|},\ \forall m_{1}\in\mathbb{N}_{0}^{n},m_{2}\in\mathbb{N}_{0}^{N},x\in U,|\xi|\geq 1;\end{split}$
for some constants $\delta>0$ and $c_{m},c_{m_{1},m_{2}}>0$.
@@
is denoted $L^{r}(U)$. By [55, Theorem 2.1, Section I.2], any $\Psi$DO in $L^{r}(U)$ is $H_{0}^{s}(U_{b})\to H_{loc}^{s-r}(U)$ continuous for any domain $U_{b}\Subset U$. The notation $\mathcal{U}_{b}\Subset\mathcal{U}$ means that the closure $\overline{\mathcal{U}_{b}}$ is compact and $\overline{\mathcal{U}_{b}}\subset\mathcal{U}$.
@@
be a nondegenerate phase function [19, Definition 11.1]. The latter property means that
@@
    The differentials $\text{d}_{(x,y,\Theta)}\Phi_{\Theta_{l}}^{\prime}$, $1\leq l\leq N$, are linearly independent on the set $\Sigma$:
    (2.10) $\Sigma:=\{(y,x,\Theta)\in V\times U\times(\mathbb{R}^{N}\setminus 0_{N}):\ \Phi_{\Theta}^{\prime}(x,y,\Theta)=0_{N}\}.$
The phase $\Phi$ determines the homogeneous canonical relation [56, Chapter VIII, Definition 5.1]
(2.11) $\begin{split}\mathcal{C}:=&\{(y,\text{d}_{y}\Phi(x,y,\Theta));(x,-\text{d}_{x}\Phi(x,y,\Theta)):(x,y,\Theta)\in\Sigma\}\\ \subset&(T^{*}V\setminus 0_{2n})\times(T^{*}U\setminus 0_{2n}).\end{split}$
$I^{r}(V\times U\times\mathcal{C})$ denotes the vector space of FIOs of order $r$ given by [19, Section 11]